4. (a) Suppose a particle P is moving in the plane so that its coordinates are given by P(x,y), where x = 4cos2t, y = 7sin2t.
(i) By finding a, b ∈ R such that x^2/a^2+y^2/ b^2 = 1, show that P is travelling on an elliptical
path.
(ii) Let L(t) be the distance from P to the origin. Obtain an expression for L(t).
(iii) How fast is the distance between P and the origin changing when t = π/8?
1
Expert's answer
2020-06-28T18:49:33-0400

This is a question of parameters so we will arrange the variables to get the desired equation of ellipse.

We will attempt all the questions sectionwise:

(i) x = 4cos2t and y = 7sin2t

Rearranging the terms

"\frac{x}{4}" = cos2t ....(i)

"\frac{y}{7}" = sin2t ....(ii)

Squaring and adding equations (i) and (ii),we get

"\frac{x^2}{16}" + "\frac{y^2}{49}" = cos22t + sin22t = 1 ; "\frac{x^2}{16}" + "\frac{y^2}{49}" = 1

Therefore comparing to standard equation of ellipse "\frac{x^2}{a^2}" + "\frac{y^2}{b^2}" = 1, we can say

a = 4 ; b =7 and equation "\frac{x^2}{16}" + "\frac{y^2}{49}" = 1 resembles the equation of ellipse. ...(answer)


(ii) L(t) is distansce between origin(0,0) and P(x,y) which is given by:

L(t) = "\sqrt{(x-0)^2 + (y-0)^2}" = "\sqrt{x^2+y^2}"

Now putting x = 4cos2t and y = 7sin2t, we get

L(t) = "\sqrt{16cos^22t+49sin^22t}" ..... (iii) .....(answer)

Therefore equation (iii) above is required equation of L(t)


(iii) To indicate how fast is distance between P and origin is changing we need to differentiate L(t) which represents the distance between them.

So differentiating equation(iii) both sides we get,

"\frac{dL(t)}{dt}" = "\frac{1}{2}" "\frac{1}{\sqrt{16cos^22t + 49sin^22t}}" ("16*2cos2t*(-2sin2t)"+ "49*2sin2t*2cos2t)"


Now rearranging and putting t="\frac{\Pi}{8}" we get

"\u200b" "\frac{dL(t)}{dt}" ="\frac{1}{2}" "\frac{1}{\sqrt{16cos^22t + 49sin^22t}}" ("98sin4t-32sin4t)"


"\frac{dL(t)}{dt}" ="\frac{1}{2}" "\frac{1}{\sqrt{16cos^2\frac{\Pi}{4} + 49sin^2\frac{\Pi}{4}}}" ("98sin\frac{\Pi}{2}-32sin\frac{\Pi}{2})"


"\frac{dL(t)}{dt}" ="\frac{1}{2}" "\frac{1}{\sqrt{16*\frac{1}{2} + 49*\frac{1}{2}}}" (98-32)


"\frac{dL(t)}{dt}" ="\frac{1}{2}" "\frac{1}{\sqrt{8+ 24.5}}" *66 = 33*"\frac{1}{\sqrt{32.5}}" = 5.788


Hence,how fast is the distance between P and the origin changing when t = "\frac{\Pi}{8}" is given by:

"\frac{dL(t)}{dt}" = 5.788 .....(answer)


Thanks and hope all doubts related are solved.